OpenStudy (anonymous):

f(x)=sqrt(x+3); find f'(1)

OpenStudy (anonymous):

find f'(x) and then plug 1 into it to get the answer

OpenStudy (anonymous):

im having trouble understanding how to find f'(x) though i dont understand the square root part of it

OpenStudy (anonymous):

x=2

OpenStudy (anonymous):

find f(1) is basically asking you to plug a "1" in everywhere you see a "x"

OpenStudy (anonymous):

f(1)= sqrt(1+3) f(1)=sqrt(4) f(1)=2

OpenStudy (anonymous):

f(1) = 2 or f'(1) = 2 ??, im lost

OpenStudy (anonymous):

f'

OpenStudy (anonymous):

i didn't see that ' there sorry :X

OpenStudy (anonymous):

i dont have 2 as an option for an answer

OpenStudy (anonymous):

hmm that's weird...what are the answers ?

OpenStudy (anonymous):

1/4, sqrt7/4, 5sqrt7/4, 5/4

OpenStudy (anonymous):

O__O

OpenStudy (anonymous):

sorry Meepi, I post it before seeing you are working on.

OpenStudy (anonymous):

f(x)=sqrt(x+3); find f'(1), can you help me with this one please. the answer appears to me to be 1/2, but it's not

OpenStudy (anonymous):

f' is the derivative, \(\Large f(x) = \sqrt{x +3}=(x + 3)^{{1 \over 2}}\) Use the chainrule + power rule to compute the derivative: \(\Large f'(x) = \frac{1}{2\sqrt{x+3}}*1=\frac{1}{2\sqrt{x+3}}\) \(\Large f'(1) = \frac{1}{2\sqrt{1 + 3}} = \frac{1}{3}\)

OpenStudy (anonymous):

whoops that \(\frac{1}{3}\) should be \(\frac{1}{4}\)

OpenStudy (anonymous):

okay 1/4 is an option so i'll take it

OpenStudy (anonymous):

still there?

OpenStudy (anonymous):

Yeah

OpenStudy (anonymous):

f(x)=sqrt(x+3); find f'(1), can you help me with this one please. the answer appears to me to be 1/2, but it's not

OpenStudy (anonymous):

Do you know how to take a derivative?

OpenStudy (anonymous):

in some cases... i know this is basic, but im lost

OpenStudy (anonymous):

i dont get it when its root or fractions

OpenStudy (anonymous):

Did you learn the chain rule?

OpenStudy (anonymous):

nope

OpenStudy (anonymous):

We can start easy, do you know the derivate of \(\sqrt{x}\)? You can use the power rule to calculate it if you write it as \(\Large x^{1 \over 2}\)

OpenStudy (anonymous):

yes i get that

OpenStudy (anonymous):

Alright so you should get: \(\Large \frac{1}{2}x^{-\frac{1}{2}}=\frac{1}{2}\frac{1}{\sqrt x}=\frac{1}{2\sqrt{x}}\) Now basically what the chain rule means is that if you have a function say \(f(x) = \sqrt{x + 3}\) You can take the derivative by pretending that what's under the root is just a single variable, as long as you multiply the result by the derivative of what it actually is. so lets just pretend x + 3 is one variable u, then according to the chain rule we can differentiate it as long as we multiply it by the derivative of u: \(f(x) = \sqrt{x + 3}\) Pretend x+3 is one variable say u: \(f(u) = \sqrt{u}, u = x + 3\) take the derivative normally, but don't forget you have to multiply by the derivative of u because you just pretend it's a single variable (and it actually isn't) \(\Large f'(x) = \frac{1}{2\sqrt u}\cdot u'\) The derivative of x + 3 is just 1 so: \(\Large f'(x) = \frac{1}{2\sqrt u}\) Now just fill in u again: \(\Large f'(x) = \frac{1}{2\sqrt{x + 3}}\) \(\Large f'(1) = \frac{1}{2\sqrt{1+3}} = \frac{1}{2\sqrt{4}} = \frac{1}{2*2}=\frac{1}{4}\)

OpenStudy (anonymous):

so what do i do in the case of a fraction like f(x)= -1/x+6 finding f'(4)?

OpenStudy (anonymous):

You use the quotient rule: \(\Large f(x) = \frac{p(x)}{q(x)}\) \(\huge f'(x) = \frac{p'(x)q(x) - p(x)q'(x)}{\left(q(x)\right)^2}\)